Ticket Number: HEA-NOC/20080527-12 Ticket Status: OPEN
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080527 15:51 UTC+1 Problem Start: 20080527 17:00 UTC+1
Ticket Update: 20080527 15:51 UTC+1 Problem End:
Site/line: CAO, BGP Routing
Bringing up second BGP session to CAO on cr2-kp
None, the primary link to the CAO is established on cr2-cwt already.
20080527 15:39 OMG New FE connection has been delivered by BT for the CAO.
I will be no shutting the BGP session and checking that all prefixes are
announced and received correctly on both ends. Will do a fail over test to
the new link to check that resilience is working properly. The new circuit
will remain under test for a week once the BGP session is established
before being accepted.
This ticket can be monitored at http://www.hea.net/tickets/20080527-12
HEAnet Limited. Registered in Ireland, No. 275301.